# Pragmatic Inference Chain (PIC) Improving LLMs’ Reasoning of Authentic Implicit Toxic Language

### v1

@@ 4 Results and Discussions
Command GPT-4o Llama-3.1 DeepSeek-v2.5 Average
Zero-shot 63.95 55.03 44.97 54.65
CoT Wei et al. (2023) 58.46 47.00 51.61 52.36
PIC one shot 69.56 51.26 55.00 58.61
PIC step instructions 76.21 68.82 64.88 69.97
PIC step instructions + three shots 74.21 53.84 71.01 66.35
For all three models, the incorporation of the PIC step instructions into the prompts significantly improves their performance. Compared to the zero-shot baseline, the PIC step instructions have brought about an increase of 12. 26% in classification accuracy with GPT, 13.79% with Llama, and 19.91% with DeepSeek.
In contrast, the highest accuracy achieved by zero-shot prompts is 63.95% with GPT. DeepSeek, which was believed to outperform the other two models in Chinese reasoning Liu et al. (2024), has the lowest accuracy, 44.97% in the pragmatic inference of implicit toxic language. More interestingly, CoT even reduces the model accuracy compared to the zero-shot baseline. GPT performance drops by 5.49% and Llama by 8.03%. DeepSeek is the only one that has a small gain of 6.64%. It may have benefited from the Chinese language that the prompts use.
@@
Furthermore, the PIC enables the models to produce more explicit and coherent reasoning processes. Figure 4 demonstrates the different inferential processes facilitated by CoT, one-shot, and step instructions. While all three prompts have led GPT to take several steps in making the inference, the PIC step instructions particularly facilitate the model to ‘notice’ more linguistic details (e.g., {CJK}UTF8gbsn "’腩’ is a pun based on the homophone of ’男’ (man)"), connect the details to common knowledge (e.g., "and also refers to “sirloins" as an ingredient"), select the knowledge that is suitable in the context (e.g., ’sirloin’ and ’cook’), and reconstruct the fundamental layer of semantic meanings (e.g., "born solely to entertain me" reconstructed as "men exist merely to provide entertainment or amusement for others"). In contrast, the inferential processes drawn on one-shot and CoT prompts tend to be unspecific and make arbitrary connections between the text and sarcasm. As a result, only the PIC step instructions are successful in identifying the implicit toxicity of this comment.
Exemplars (shots) vary their effects across different models and, at times, add little to the model improvement. Unlike prior studies that identified improvements from in-context learning of concrete shots (e.g., Ma et al., 2023; Nachane et al., 2024), one-shot alone in our study only increases the GPT performance by 5.61% and DeepSeek by 10.03%, while decreasing Llama performance by 3.77%. We should emphasize that the increase/decrease is compared to the zero-shot baselines. The accuracy achieved by one-shot prompting consistently falls behind the PIC step instructions. Including more shots to the step instructions, i.e., step instructions + 3 shots, also reduces the performance of GPT by 2% and Llama by 14.98%, but increases the DeepSeek performance by 6.13%, compared to PIC step instructions only.
@@
The result may find both similarities and differences with humans’ ability to make pragmatic inferences. Previous studies of cognitive psychology studies have found that humans guide their pragmatic inference by abstract ’schemata’ – generalized sets of rules defined in relation to classes of goals Cheng and Holyoak (1985); Mazzone (2011), instead of concrete examples. Our PIC step instructions may be analogous to the rules. However, humans extract the rules from concrete and discursive exemplars, such as repeated social experiences of how certain speeches are interpreted in context Ochs (1988). LLMs appear to learn the rules directly without the need for concrete individual demonstrations.
Overall, the results show that LLMs, regardless of whether they advance in logical reasoning or not, need PIC assistance to achieve a higher performance in reasoning authentic implicit toxic language. CoT, which has been efficient in symbolic, arithmetic, and commonsense reasoning, does not perform better than directly prompting LLMs in the task. The findings suggest that previous neurological evidence of the difference between logical reasoning and pragmatic inferences Prado et al. (2015); Spotorno et al. (2015) may also apply to artificial neural networks.
@@
How different types of reasoning tasks result in different neuron activities in artificial neural networks goes beyond the scope of the current study, but we may be able to provide some interdisciplinary explanations for the efficiency of PIC. Besides the relevance theory revised in Section 2, the Noticing Hypothesis proposed by Schmidt (1990) suggests that conscious pick-up of language input is necessary for human learning of language meanings. Albeit whether LLMs are conscious is controversial, the first step of the PIC has indeed prompted the LLMs to pick up more linguistic input explicitly. This may be explained by a changing weight in their attention mechanism, which, however, needs further investigation. Chen and Lee (2021) and Chen and Brown (2024) experimentally evidence that humans build their understanding of context-specific meanings off the back of stereotypical meanings of a language. Therefore, the second step of the PIC, which requires the LLMs to explain the literal meaning of the comment, could have provided a foundation for their context-specific understanding of implicit toxicity. Lastly, the third step asks the LLMs to compare the meanings of the comment against gender equality principles, namely, bringing up the existing requirements for controlled text generation Liang et al. (2024). The potential contributions of each step can thus explain the efficiency of PIC over other prompting methods that could not entail them.
@@ 5 Related work
There has been a surge of research on improving the reasoning capability of LLMs, following the design of CoT Buhnila et al. (2024); Fang et al. (2025); Huang et al. (2025); Konya et al. (2024); Lin et al. (2024); Niu et al. (2024); Pan et al. (2025). While the majority concentrates on logical reasoning, we have also witnessed an increasing interest in rule-based reasoning Servantez et al. (2024) and reasoning through theory-of-mind Lin et al. (2024). For example, Servantez et al. (2024) was inspired by the IRAC framework (Issue, Rule, Application, and Conclusion) developed by lawyers and formulated instructive reasoning steps to improve LLMs’ accuracy in making legal decisions. Interestingly, in legal tasks, Blair-Stanek et al. (2023) also found that exemplars in prompting did not help improve LLM performance. Servantez et al. emphasized that their rule-based Chain of Logic provided LLMs with some freedom, that is, let the models “decide how many rule elements exist, the text span of each element and the logical relationships between them” (p.2722). The current PIC step instructions substantiate the role of such freedom, as it also leaves the decisions to LLMs to identify the linguistic stimuli to be ‘noticed’, the relevance between the stimuli, the context and common knowledge, and the literal meanings expressed.
Recent studies have gone beyond the grammatical accuracy and semantic coherence that LLMs seem to have achieved, and started paying more attention to their pragmatic capability. Concerning pragmatic inference, Qiu et al (2023) found the early version of ChatGPT almost unable to interpret scalar implicatures. Hu et al (2023), Ruis et al. (2023), and Barattieri Di San Pietro et al. (2023) all identified LLM’s difficulty in comprehending humour and irony. Sravanthi et al (2024) highlighted LLMs’ shortcomings in understanding pragmatic presuppositions – a preparatory stage for pragmatic inference. Despite the many pragmatic issues identified, systematic solutions have been scarce. The PIC proposed by the current study might offer one of the first systematic solutions for complex pragmatic inferential tasks in general, not restricted to the reasoning of implicit toxic language. It demonstrates that the unsatisfactory performance of LLMs in pragmatic tasks can be improved by in-context learning.
@@ 6 Conclusion
This study proposes a new in-context learning method, the Pragmatic Inference Chain (PIC), drawn on findings from cognitive science and linguistics. It also presents a newly established authentic implicit toxic dataset that requires intensive pragmatic inferences. It tests the PIC on three LLMs. The findings reveal that the PIC significantly improves their success rate of identifying implicit toxic language, compared to both zero-shot prompting and CoT. The method also enables the LLMs to move from unspecified stepped inferences to explicit and coherent inference processes. The design of the PIC may apply to other pragmatic inferential tasks, such as metaphors and humour comprehension, where LLMs are found deficient. It also helps LLMs address real-world challenges in handling the creative range of implicit toxic language use.
@@ 7 Limitations
While the PIC step instructions are found effective and exemplars add little to the result, we also observe that even one-shot prompt has led the LLM to pick up some linguistic details that are not found with CoT (see Figure 4). It thus raises the question of whether providing more shots than the current three would bring a noticeable increase in the accuracy of understanding implicit toxic language. Additionally, LLMs can now be fine-tuned by machine-generated PIC to improve further in making pragmatic inferences. Previously, the relevance-theoretical inferential procedures relied on manual production. With the proposed PIC, distillation becomes possible. However, caution needs to be paid to the machine-generated PIC, as it may not be as felicitous as human-provided ones. That is, some machine-generated PICs have not fully explained all linguistic stimuli or the literal meanings that are relevant to the pragmatic understanding, but still reached a correct conclusion (see Appendix E). How the partially completed inference processes affect fine-tuning needs further investigation.
@@ Appendix E Partial inferential process facilitated by PIC step instructions
1. The comment does not use metaphors or words with special meanings. The term "mama’s boy" generally refers to men who are overly dependent on their mothers in daily life, while "envy" expresses admiration or desire for a certain state or trait.
2. The literal meaning of the comment is: "The reason ’mama’s boys’ are looked down upon is that their mothers are strong and difficult to deal with, which is why they are discredited. In reality, many people envy ’mama’s boys.’"
3. The statement "A mama’s boy means his mother is strong and hard to take advantage of, so people try to discredit him" may imply a gender stereotype—specifically, that a strong mother leads to her son being devalued. This perspective overlooks individual independence and contradicts the principle of gender equality.

### v2

@@ 4.1 The effectiveness of PIC
For the four non-reasoning models, the PIC step instructions have significantly improved their performance.. Compared to the zero-shot baseline, the PIC step instructions alone bring about an increase of 12. 26% in the classification accuracy with GPT, 13.79% with Llama, 19.91% with DeepSeek-v2.5, and 19.14% with DeepSeek-v3. Adding a rule-based prompt to it, namely, the PIC step instructions + rule, gives a further small gain of 1% - 4.5%.
The rule-based prompt is also the only one of the five baseline methods that consistently improves the models’ performance in the current task. While the finding indicates the effectiveness of the safety principles implemented in the models, the improvements that they lead to are barely half of those of the PIC step instructions. In other words, PIC step instructions are noticeably more effective in the implicit toxicity identification, while not being more complicate to design or to apply than the safety principles.
Compared to the non-reasoning models, QwQ32b – a reasoning model that is comparable to DeepSeek-R1 in mathematical and coding tasks – shows a completely insensitivity to any of the prompts. Its success rate fluctuates only above and below the zero-shot baseline and has never been above chance. It thus appears that QwQ32b’s high performance in logical reasoning is achieved at some cost to its capability of pragmatic inference. It is unclear whether enhancing the logical reasoning ability of an LLM would reduce its capability of doing non-demonstrative reasoning. However, we do observe some collateral evidence, for example, adding CoT results in worse performance of GPT-4o and Llama in the current toxicity inference compared to their zero-shot baselines.
@@
4.2 The ’mavericks’
Although PIC step instructions improved the performance of non-reasoning models unanimously, the models demonstrate several interesting patterns with other types of prompts. For example, Llama-3.1-70B-Instruct yields a reversed performance in shot-involved prompts. It increases its performance in three-shot baseline prompt while all the other non-reasoning models decrease, and it decreases over the PIC shots while all the others increase. Recall that the difference between normal shots and PIC shots was whether they involved the inferential processes. Therefore, it seems that Llama learns the pragmatic inference better from the labeling patterns, but not from the concrete examples of the inferential process.
Similarly, the two DeepSeek models improve their success rate with CoT, when the others decrease. As a trick to improve LLMs’ capability of logical reasoning, CoT has previously been found to be not effective in nonlogical reasoning Sprague et al. (2024). This is in line with our findings on GPT and Llama. However, DeepSeek’s improvement over CoT prompts in the current task suggests another possibility. That is, CoT as an in-context learning method might not work in pragmatic inference, but after it has been embedded as part of reinforcement learning, such as post-training of DeepSeek models DeepSeek-AI et al. (2025), the prompt may trigger the models to assign different weights to their parameters and therefore becomes effective in pragmatic inference. Our arguments are partly corroborated by Chua and Evans2025 who find that non-reasoning models fine-tuned by the distillation of CoT from DeepSeek-R1 exhibit similar reasoning-like behaviours.
4.3 The interdisciplinary explanations for prompt effectiveness
Across the prompts, exemplars (shots) in general add little to the model improvement. Unlike previous studies that identified improvements from in-context learning of concrete shots (e.g., Ma et al., 2023; Nachane et al., 2024), both baseline shots and PIC shots either reduce the model performance compared to prompts without them or only provide a marginal gain.
@@ 4.3 The interdisciplinary explanations for prompt effectiveness
The result shows both similarities and differences with humans’ ability to make pragmatic inferences. Previous studies of cognitive psychology have found that humans guide their pragmatic inference by abstract ’schemata’ – generalized sets of rules defined in relation to classes of goals Cheng and Holyoak (1985); Mazzone (2011), instead of concrete examples. Our PIC step instructions may be analogous to the schemata. However, humans extract their schemata from concrete and discursive exemplars, such as repeated social experiences of how thank you is interpreted as polite in context Ochs (1988). LLMs appear to learn the schemata from step instructions directly without the need for concrete individual demonstrations.
Learning the PIC step instructions also enables the models to produce more explicit and coherent reasoning processes. Figure 4 demonstrates the different inferential processes facilitated by CoT, one-shot, and step instructions. While all three prompts have led GPT to take several steps in making the inference, the PIC step instructions particularly facilitate the model to ‘notice’ more linguistic details (e.g., ”’腩’ is a pun based on the homophone of ’男’ (man)”), connect the details to common knowledge (e.g., ”and also refers to “sirloins” as an ingredient”), select the knowledge that is suitable in the context (e.g., ’sirloin’ and ’cook’), and reconstruct the fundamental layer of semantic meanings (e.g., ”born solely to entertain me” reconstructed as ”men exist merely to provide entertainment or amusement for others”). In contrast, the inferential processes drawn on one-shot and CoT prompts tend to be unspecific and make arbitrary connections between the text and sarcasm. As a result, only the PIC step instructions are successful in identifying the implicit toxicity of this comment.
@@
The efficiency of PIC instruction steps may find some interdisciplinary explanations from linguistics and cognitive science. Besides the relevance theory revised in Section 2, the Noticing Hypothesis proposed by Schmidt (1990) suggests that conscious pick-up of language input is necessary for human learning of language meanings. Albeit whether LLMs are conscious is controversial, the first step of the PIC has indeed prompted the LLMs to pick up more linguistic input explicitly. This may be explained by a changing weight in their attention mechanism, which is worth further investigation. Chen and Lee (2021) and Chen and Brown (2024) experimentally evidence that humans build their understanding of context-specific meanings off the back of conventional meanings of a language. Therefore, the second step of the PIC, which requires the LLMs to explain the literal meaning of the comment, could have provided a foundation for their context-specific understanding of implicit toxicity. Finally, the third step asks the LLMs to compare the meanings of the comment against gender equality principles, namely, bringing up the existing requirements for controlled text generation Liang et al. (2024). The potential contributions of each step may have boosted the success rate of PIC over other prompting methods that could not entail them.
We should note that PIC prompts are not always effective. There are approximately 7.5% of the data where all five models failed to identify the (non)toxicity. Scrutinizing these failed cases shows that they often contain complex perspective-taking practices when being toxic, e.g., males taking on the viewpoint of females to be sarcastic about female behaviours. Since 2023, a very small number of studies have realized the power of perspective-taking in diminishing toxicity and enhancing LLMs’ reasoning Just et al. (2024); Xu et al. (2024); Wilf et al. (2023). They derived their prompt design from findings in social psychology or cognitive science. Perspective-taking has also been studied as ‘footing’ and ‘stance’ in pragmatics Butler (2007); Goffman (1981). Leveraging their insight, future studies are encouraged to explore the potential of adding a step on perspective discernment into the PIC design.
@@ 5 Related work
Thus far, LLMs’ capability of doing logical reasoning has been one of the rapidly growing topics in LLM research. We have witnessed the surge of different CoT designs Buhnila et al. (2024); Fang et al. (2025); Huang et al. (2025); Konya et al. (2024); Lin et al. (2024); Niu et al. (2024); Pan et al. (2025) and the development of various reasoning models. This paper, however, demonstrates that logical reasoning is only one piece of the puzzle in advancing LLMs’ reasoning ability. Other reasoning abilities, such as pragmatic inference, are equally crucial to the LLMs’ performance, but has been much more underexplored. Noticed the research gap, several studies have explored rule-based reasoning Servantez et al. (2024) and reasoning through theory-of-mind Lin et al. (2024). For example, Servantez et al. (2024) was inspired by the IRAC framework (Issue, Rule, Application, and Conclusion) developed by lawyers and formulated instructive reasoning steps to improve LLMs’ accuracy in making legal decisions. Interestingly, in legal tasks, Blair-Stanek et al. (2023) also found that exemplars in prompting did not help improve LLM performance. Servantez et al. emphasized that their rule-based Chain of Logic provided LLMs with some freedom, that is, let the models “decide how many rule elements exist, the text span of each element and the logical relationships between them” (p.2722). The current PIC step instructions substantiate the role of such freedom, as it also leaves the decisions to LLMs to identify the linguistic stimuli to be ‘noticed’, the relevance between the stimuli, the context and common knowledge, and the literal meanings expressed.
Recent studies have also gone beyond the grammatical accuracy and semantic coherence of LLM generation, and started paying more attention to their pragmatic capability. Concerning pragmatic inference, Qiu et al (2023) found the early version of ChatGPT almost unable to interpret scalar implicatures. Hu et al (2023), Ruis et al. (2023), and Barattieri Di San Pietro et al. (2023) all identified LLM’s difficulty in comprehending humour and irony. Sravanthi et al (2024) highlighted LLMs’ shortcomings in understanding pragmatic presuppositions – a preparatory stage for pragmatic inference. Despite the many pragmatic issues identified, systematic solutions have been scarce. The PIC proposed by the current study might offer one of the first systematic solutions for complex pragmatic inferential tasks in general, not restricted to the reasoning of implicit toxic language. It demonstrates that the unsatisfactory performance of LLMs in pragmatic tasks can be improved by in-context learning.
@@ 6 Conclusion
This study proposes a new in-context learning method, the Pragmatic Inference Chain (PIC), drawn on findings from cognitive science and linguistics. It also presents a newly established authentic implicit toxic dataset that requires intensive pragmatic inferences. It tests varied PIC designs, together with five baseline prompts, on five LLMs. The findings reveal that the PIC significantly improves the models’ success rate of identifying implicit toxic language, compared to all baselines. The method also enables the LLMs to move from unspecified stepped inferences to explicit and coherent inference processes. The design of the PIC may apply to other pragmatic inferential tasks, such as metaphors and humour comprehension, where LLMs are found deficient. It also helps LLMs address real-world challenges in handling the creative range of implicit toxic language use.
@@ 7 Limitations
While the PIC step instructions are found effective and exemplars add little to the result, we also observe that even one-shot PIC prompt has led the LLM to pick up some linguistic details that are not found with CoT (see Figure 4). It thus raises the question of whether providing more shots of PIC than the current three would bring a noticeable increase in the accuracy of understanding implicit toxic language. Additionally, LLMs can now be fine-tuned by machine-generated PIC to improve further in making pragmatic inferences. Previously, the relevance-theoretical inferential procedures relied on manual production. With the proposed PIC step instructions, distillation becomes possible. However, caution needs to be paid to the machine-generated PIC, as it may not be as felicitous as human-provided ones. That is, some machine-generated PICs have not fully explained all linguistic stimuli or the literal meanings that are relevant to the pragmatic understanding, but still reached a correct conclusion (see Appendix E). How the partially completed inference processes affect fine-tuning needs further investigation.
@@ Appendix E Partial inferential process facilitated by PIC step instructions
1. The comment does not use metaphors or words with special meanings. The term ”mama’s boy” generally refers to men who are overly dependent on their mothers in daily life, while ”envy” expresses admiration or desire for a certain state or trait.
2. The literal meaning of the comment is: ”The reason ’mama’s boys’ are looked down upon is that their mothers are strong and difficult to deal with, which is why they are discredited. In reality, many people envy ’mama’s boys.’”
3. The statement ”A mama’s boy means his mother is strong and hard to take advantage of, so people try to discredit him” may imply a gender stereotype—specifically, that a strong mother leads to her son being devalued. This perspective overlooks individual independence and contradicts the principle of gender equality.